# On Erlang mixture approximations for differential equations with distributed time delays

### v1

@@ 4.1 Convergence
Let $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be a delta family parametrized by $a\in\mathbb{R}_{>0}$ which satisfies the following conditions.
@@
    For all $\gamma\in\mathbb{R}_{>0}$ and for all $s\in\mathbb{R}_{\geq 0}$ satisfying $\gamma\leq|t-s|<\infty$, $\delta_{a}(t,s)\rightarrow 0$ uniformly as $a\rightarrow\infty$ for all $t\in\mathbb{R}_{\geq 0}$.
@@
    $\delta_{a}(t,s)\geq 0$ for all $t,s\in\mathbb{R}_{\geq 0}$ and $a\in\mathbb{R}_{>0}$.
@@
(46) $=\underbrace{\int_{[t-\gamma]^{+}}^{t+\gamma}\delta_{a}(t,s)\alpha(s)\,\mathrm{d}s}_{\mathcal{I}_{1}(t)}+\underbrace{\int_{0}^{[t-\gamma]^{+}}\delta_{a}(t,s)\alpha(s)\,\mathrm{d}s+\int_{t+\gamma}^{\infty}\delta_{a}(t,s)\alpha(s)\,\mathrm{d}s}_{\mathcal{I}_{2}(t)}.$
@@
(47) $\mathcal{I}_{1}(t)$ $=\underbrace{\alpha(t)\int_{[t-\gamma]^{+}}^{t+\gamma}\delta_{a}(t,s)\,\mathrm{d}s}_{\mathcal{I}_{3}(t)}+\underbrace{\int_{[t-\gamma]^{+}}^{t+\gamma}\delta_{a}(t,s)(\alpha(s)-\alpha(t))\,\mathrm{d}s}_{\mathcal{I}_{4}(t)}.$
@@
Let $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be an Erlang mixture kernel of infinite order with rate parameter $a\in\mathbb{R}_{>0}$,
(53) $\hat{\alpha}(t)$ $=\sum_{m=0}^{\infty}c_{m}\ell_{m}(t),$
where the coefficients are given by
(54) $c_{m}$ $=\int_{s_{m}}^{s_{m+1}}\alpha(s)\,\mathrm{d}s,$ $m$ $=0,1,\ldots,$
for $s_{m}=m\Delta s\in\mathbb{R}_{\geq 0}$ and $\Delta s=1/a\in\mathbb{R}_{>0}$. Furthermore, let $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be a regular kernel. Then, $\hat{\alpha}$ converges pointwise to $\alpha$, i.e.,
(55) $\hat{\alpha}(t)\rightarrow\alpha(t)$
as $a\rightarrow\infty$ for all $t\in\mathbb{R}_{\geq 0}$.
Proof 4.25.
First, we prove that the Erlang mixture delta family $\delta_{a}$ satisfies the three properties of Theorem 4.22. The first property follows directly from Lemma 4.20. The second property follows from Lemma 4.10 and the monotonicity properties in Lemma 4.12, i.e., the convergence is uniform outside of an interval around $t$ because $\delta_{a}$ is monotonically non-decreasing and non-increasing for values of $s$ lower and higher than $t$, respectively. Finally, it is straightforward to verify the third property from Definition 4.1 and 4.3 because $\ell_{m}$ is the product of a monomial and an exponential function with non-negative arguments. Next, it follows from the definition of the Erlang mixture delta family, $\delta_{a}$, that the approximation in (44) is an Erlang mixture kernel of infinite order,
(56) $\hat{\alpha}(t)$ $=\int_{0}^{\infty}\delta_{a}(t,s)\alpha(s)\,\mathrm{d}s=\sum_{m=0}^{\infty}\ell_{m}(t)\int_{s_{m}}^{s_{m+1}}\alpha(s)\,\mathrm{d}s,$
and that the coefficients are given by (54).
Corollary 4.26.
The coefficients defined by (54) sum to one, i.e.,
(57) $\sum_{m=0}^{\infty}c_{m}$ $=1.$
@@
This is a straightforward consequence of the fact that, by assumption, the integral of $\alpha$ is one. We write out the sum of the coefficients:
Remark 4.28.
To the best of the author’s knowledge, the result in Theorem 4.24 is novel, i.e., it has not been proven that the probability density function of an Erlang mixture distribution of infinite order converges pointwise to that of any non-negative random variable with a density function that satisfies the regularity conditions in Definition 3.1. It has been proven [11, 49] that the cumulative distribution function of such an Erlang mixture distribution converges to that of any non-negative random variable (i.e., convergence in distribution has been proven). Furthermore, it is well-known that, according to Scheffé’s lemma [45], the convergence of probability density functions implies convergence of cumulative distribution functions. However, the converse is only true if the probability density function is bounded and equicontinuous [7, 48]. Alternatively, the converse is also true if the probability density function converges uniformly [44]. However, it is outside the scope of the present work to further investigate these alternative proofs.
@@ 5 Approximate system of ordinary differential equations
In this section, we use the Erlang mixture approximation presented in Section 4 to approximate the multivariate kernel in the DDEs presented in Section 3, and we use the LCT to transform the resulting approximate system to ODEs. Furthermore, we describe the steady state equations for the approximate system of ODEs together with a stability criterion. Finally, we present a comparison of the stability criteria for the approximate and original system.
@@
where $\hat{Z}:\mathbb{R}\rightarrow\mathbb{R}^{M+n_{z}}$ contains auxiliary memory states, and $M=\sum_{i=1}^{n_{z}}M_{i}$. Furthermore, the matrices $A\in\mathbb{R}^{M+n_{z}\times M+n_{z}}$, $B\in\mathbb{R}_{\geq 0}^{M+n_{z}\times n_{z}}$, $C\in[0,1]^{n_{z}\times M+n_{z}}$ are block-diagonal:
@@
(68) $\left(A^{(i)}\right)^{-1}$ $=\frac{-1}{a}\begin{bmatrix}1\\ 1&1\\ \vdots&&\ddots\\ 1&1&\cdots&1\end{bmatrix},$ $c^{(i)}\left(A^{(i)}\right)^{-1}b^{(i)}$ $=-\sum_{m=0}^{M_{i}}c_{m,i}=-1.$
Consequently, $-CA^{-1}B=I$ is an identity matrix, and
@@
where $F$, $G$, and $H$ are the Jacobian matrices given by (8). Furthermore, the eigenvalues can be partitioned into two subsets. One of the subsets will contain $M_{i}$ eigenvalues equal to $-a_{i}$ for $i=1,\ldots,n_{z}$, and the other will contain eigenvalues that satisfy the reduced characteristic equation
(71) $\det(F-\lambda I+GQ(\lambda)H)$ $=0,$ $Q_{ij}(\lambda)$ =$=\begin{cases}\sum\limits_{m=0}^{M_{i}}c_{m,i}\left(\frac{a_{i}}{a_{i}+\lambda}\right)^{m+1},&i=j,\\ 0,&i\neq j,\end{cases}$
where $Q:\mathbb{C}\rightarrow\mathbb{C}^{n_{z}\times n_{z}}$.
@@
where we have used Schur’s determinant formula (see, e.g., [55]). Next, we exploit the block-diagonal structure of $A$:
@@
and $\lambda=-a_{i}$ is an eigenvalue with algebraic multiplicity $M_{i}$ for $i=1,\ldots,n_{z}$. As the rate parameters, $a_{i}$, are positive, $A-\lambda I$ is invertible. Next, we consider the diagonal matrix
@@
Proof 5.6 (Alternative proof).
For completeness, we demonstrate that the result can also be obtained from the equivalence between (59) and (61) and the stability criterion in Lemma 3.4. First, we write out the following integral for an Erlang kernel of order $m$ with rate parameter $a_{i}$:
@@
and the characteristic equation (7) becomes
(83) $\det(F-\lambda I+GQ(\lambda)H)$ $=0.$
Next, we analyze the relation between the stability criterion from Theorem 5.4 for the approximate system of ODEs and the one from Theorem 3.4 for the original system of DDEs. The analysis is based on the piecewise constant approximation of decaying exponentials described in Lemma 5.7 and illustrated in Figure 3.
Figure 3: Left: The piecewise constant approximation (84) from Lemma 5.7 of a decaying exponential function, $e^{-\lambda s}$ for $\lambda=1/2$ and different values of $a$. Right: The absolute error of the approximation.
Lemma 5.7.
The approximation
(84) $e^{-\lambda s}$ $\approx\left(\frac{a}{a+\lambda}\right)^{m+1},$ $s$ $\in[s_{m},s_{m+1}),$ $s_{m}$ $=m\Delta s,$ $\Delta s$ $=1/a,$
converges as $a\rightarrow\infty$ for all $s,\lambda\in\mathbb{R}_{\geq 0}$.
Proof 5.8.
First, we note that
(85) $\frac{a}{a+\lambda}$ $=1-\frac{\lambda}{a+\lambda},$ $\left(1-\frac{\lambda}{a+\lambda}\right)^{m+1}$ $\approx e^{-(m+1)\frac{\lambda}{a+\lambda}},$
where the approximation is obtained by truncating the sum in (1) after the two first terms. It is valid for small arguments, and $\lambda/(a+\lambda)\rightarrow 0$ as $a\rightarrow\infty$. Furthermore, $m+1=s_{m+1}/\Delta s\approx as$ such that
(86) $e^{-(m+1)\frac{\lambda}{a+\lambda}}$ $\approx e^{-\frac{a}{a+\lambda}\lambda s}\rightarrow e^{-\lambda s}$
as $a\rightarrow\infty$. For $\lambda=0$, the approximation is exact for all nonzero values of $a$.
Theorem 5.9.
For infinite-order Erlang mixture approximations of the elements of the kernel, $\alpha$, as described in Theorem 4.24, the reduced characteristic function in (71) converges to the characteristic function in (7) as $a_{i}\rightarrow\infty$ for $i=1,\ldots,n_{z}$.
Proof 5.10.
First, we approximate the exponential function in (7) by
(87) $e^{-\lambda s}$ $\approx\left(\frac{a_{i}}{a_{i}+\lambda}\right)^{m+1},$ $s$ $\in[s_{m},s_{m+1}),$ $s_{m}$ $=m\Delta s,$ $\Delta s$ $=1/a_{i},$
for the $i$’th kernel. Next, we consider the integral in (7) for the $i$’th element of $\alpha$ and substitute the approximation:
(88) $\int_{0}e^{-\lambda s}\alpha_{i}(s)\,\mathrm{d}s$ $\approx\sum_{m=0}^{\infty}\left(\frac{a_{i}}{a_{i}+\lambda}\right)^{m+1}\int_{s_{m}}^{s_{m+1}}\alpha_{i}(s)\,\mathrm{d}s=\sum_{m=0}^{\infty}c_{m,i}\left(\frac{a_{i}}{a_{i}+\lambda}\right)^{m+1}=Q_{ii}(\lambda).$
The coefficients were defined in (54) in Theorem 4.24. Consequently, as the approximation of the exponential function in (87) converges as $a_{i}\rightarrow\infty$ for $i=1,\ldots,n_{z}$,
(89) $\det(F-\lambda I+GQ(\lambda)H)\rightarrow\det\left(F-\lambda I+G\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\alpha(s)\,\mathrm{d}sH\right),$
as well, which concludes the proof.
@@
Proof 5.12.
As the rate parameters, $a_{i}$, are positive for $i=1,\ldots,n_{z}$, an eigenvalue of zero must be a root of the reduced characteristic equation in (71), i.e., it is not an eigenvalue of $A$. Furthermore,

### v2

@@ 4.2 Pointwise convergence
Let $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be a delta family parametrized by $a\in\mathbb{R}_{>0}$ which satisfies the following conditions for any given $t\in\mathbb{R}_{\geq 0}$.
@@
    For any $\gamma\in\mathbb{R}_{>0}$, $\delta_{a}(t,s)\rightarrow 0$ uniformly for $s\in\mathbb{R}_{\geq 0}$ satisfying $\gamma\leq|t-s|<\infty$ as $a\rightarrow\infty$.
@@
    For all $s\in\mathbb{R}_{\geq 0}$ and $a\in\mathbb{R}_{>0}$, the delta family is non-negative, i.e., $\delta_{a}(t,s)\geq 0$.
@@
(51) $=\underbrace{\int_{[t-\gamma]^{+}}^{t+\gamma}\delta_{a}(t,s)\alpha(s)\,\mathrm{d}s}_{\mathcal{I}_{1}(t)}+\underbrace{\int_{0}^{[t-\gamma]^{+}}\delta_{a}(t,s)\alpha(s)\,\mathrm{d}s+\int_{t+\gamma}^{\infty}\delta_{a}(t,s)\alpha(s)\,\mathrm{d}s}_{\mathcal{I}_{2}(t)}\hskip-2.0pt.$
@@
(52) $\mathcal{I}_{1}(t)$ $=\underbrace{\alpha(t)\int_{[t-\gamma]^{+}}^{t+\gamma}\delta_{a}(t,s)\,\mathrm{d}s}_{\mathcal{I}_{3}(t)}+\underbrace{\int_{[t-\gamma]^{+}}^{t+\gamma}\delta_{a}(t,s)(\alpha(s)-\alpha(t))\,\mathrm{d}s}_{\mathcal{I}_{4}(t)}\hskip-2.0pt.$
@@
Let $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be an Erlang mixture kernel of infinite order with coefficients defined by (17) and rate parameter $a\in\mathbb{R}_{>0}$, and let $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be a regular kernel. Then, $\hat{\alpha}$ converges pointwise to $\alpha$ as $a\rightarrow\infty$ for all $t\in\mathbb{R}_{\geq 0}$.
@@
First, we prove that the Erlang mixture delta family $\delta_{a}$ satisfies the three properties of Theorem 4.24. The first property follows directly from Lemma 4.22. The second property follows from Lemma 4.12 and the monotonicity properties in Lemma 4.14, i.e., the convergence to zero is uniform outside of an interval around $t$ because $\delta_{a}$ is monotonically non-decreasing and non-increasing for values of $s$ lower and higher than $t$, respectively. Finally, it is straightforward to verify the third property from Definition 4.1 and 4.3 because $\ell_{m}$ is the product of a monomial and an exponential function with non-negative arguments.
4.3 Convergence of the integral
In this section, we briefly state a previous result presented by Tijms [51] on the convergence of the integral of an infinite-order Erlang mixture approximation, $\hat{\alpha}$, to that of any regular kernel, $\alpha$. We will use this result to prove properties of the Erlang mixture approximation in Section 4.4 and to prove the convergence of the integral of the absolute error in Section 4.5.
Theorem 4.28 (Tijms [51, Thm. 2.9.1]).
Let $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be an infinite-order Erlang mixture kernel with rate parameter $a\in\mathbb{R}_{>0}$ and coefficients given by (17), and let $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be a regular kernel. Then, for all $t\in\mathbb{R}_{\geq 0}$, the integral of $\hat{\alpha}$ converges pointwise to that of $\alpha$, i.e.,
(58) $\int_{0}^{t}\hat{\alpha}(s)\,\mathrm{d}s\rightarrow\int_{0}^{t}\alpha(s)\,\mathrm{d}s,$
as $a\rightarrow\infty$.
4.4 Properties of Erlang mixture kernels
In this section, we prove four properties of infinite-order Erlang mixture kernels with coefficients given by (17) that will be used either to prove convergence of the integral of the absolute error in Section 4.5 or to prove a relation between the stability analyses of the original system in Section 3 and the approximate system presented in Section 5.
Lemma 4.29.
For given $a\in\mathbb{R}_{>0}$, the infinite-order Erlang mixture kernel $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ obtained with the coefficients (17) is bounded if the kernel $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is bounded. Specifically, if $\alpha(t)\leq K$ for all $t\in\mathbb{R}_{\geq 0}$, then $\hat{\alpha}(t)\leq K$ as well.
Proof 4.30.
The boundedness follows directly from the substitution of the upper bound on $\alpha$ into the expression (17) for the coefficients:
(59) $\hat{\alpha}(t)$ $=\sum_{m=0}^{\infty}\ell_{m}(t)\int_{s_{m}}^{s_{m+1}}\alpha(s)\,\mathrm{d}s\leq\sum_{m=0}^{\infty}\ell_{m}(t)\int_{s_{m}}^{s_{m+1}}K\,\mathrm{d}s=K\sum_{m=0}^{\infty}\ell_{m}(t)\Delta s$
$=K\sum_{m=0}^{\infty}\frac{(at)^{m}}{m!}ae^{-at}\frac{1}{a}=K\sum_{m=0}^{\infty}\frac{(at)^{m}}{m!}e^{-at}=Ke^{at}e^{-at}=K,$
where we have used the identity (1).
Lemma 4.31.
For given $a\in\mathbb{R}_{>0}$, the infinite-order Erlang mixture kernel $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ obtained with the coefficients (17) is exponentially bounded if the kernel $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is exponentially bounded. Specifically, if $\alpha(t)\leq Le^{-\rho t}$ for all $t\in\mathbb{R}_{\geq 0}$ and $a\geq\rho/\omega$ for some $\omega\in\mathbb{R}_{>0}$, then $\hat{\alpha}(t)\leq Le^{-\bar{\rho}t}$ where $\bar{\rho}=\rho(1-e^{-\omega})/\omega\in\mathbb{R}_{>0}$. Furthermore, $0\leq\bar{\rho}<\rho$ and $\bar{\rho}\rightarrow\rho$ as $\omega\rightarrow 0$ (corresponding to $a\rightarrow\infty$).
Proof 4.32.
First, we bound the following exponential [^1] The convexity of decaying exponential functions implies that $e^{-x}\leq 1-\frac{1-e^{-\omega}}{\omega}x$ for $x\in[0,\omega]$. for $\rho/a\leq\omega$:
(60) $e^{-\rho s_{m}}$ $=e^{-\rho m/a}=\left(e^{-\rho/a}\right)^{m}\leq\left(1-\frac{1-e^{-\omega}}{\omega}\frac{\rho}{a}\right)^{m}=\left(1-\frac{\bar{\rho}}{a}\right)^{m}=\left(\frac{a-\bar{\rho}}{a}\right)^{m}.$
Next, we substitute the exponential bound in the expression (17) for the coefficients:
(61) $c_{m}$ $=\int_{s_{m}}^{s_{m+1}}\alpha(s)\,\mathrm{d}s\leq L\int_{s_{m}}^{s_{m+1}}e^{-\rho s}\,\mathrm{d}s=-\frac{L}{\rho}(e^{-\rho s_{m+1}}-e^{-\rho s_{m}})$
$=\frac{L}{\rho}e^{-\rho s_{m}}\left(1-e^{-\rho(s_{m+1}-s_{m})}\right)=\frac{L}{\rho}e^{-\rho s_{m}}(1-e^{-\rho\Delta s})\leq Le^{-\rho s_{m}}\Delta s=Le^{-\rho s_{m}}\frac{1}{a}$
$\leq L\left(\frac{a-\bar{\rho}}{a}\right)^{m}\frac{1}{a}.$
Here, we have used the bound $1-e^{-\rho\Delta s}\leq\rho\Delta s$. Next, we substitute this bound into the Erlang mixture kernel:
(62) $\hat{\alpha}(t)$ $=\sum_{m=0}^{\infty}c_{m}\ell_{m}(t)\leq\sum_{m=0}^{\infty}L\left(\frac{a-\bar{\rho}}{a}\right)^{m}\frac{1}{a}\frac{(at)^{m}}{m!}ae^{-at}=L\sum_{m=0}^{\infty}\frac{((a-\bar{\rho})t)^{m}}{m!}e^{-at}$
$=Le^{-\bar{\rho}t}\sum_{m=0}^{\infty}\frac{((a-\bar{\rho})t)^{m}}{m!}e^{-(a-\bar{\rho})t}=Le^{-\bar{\rho}t}e^{(a-\bar{\rho})t}e^{-(a-\bar{\rho})t}=Le^{-\bar{\rho}t}.$
Again, we have used the identity (1). The bound $1-e^{-\omega}\leq\omega$ provides the bound $\bar{\rho}\leq\rho$, and using the identity (1), we show that $\bar{\rho}\rightarrow\rho$ as $\omega\rightarrow 0$:
(63) $\bar{\rho}$ $=\rho\frac{1-e^{-\omega}}{\omega}=\rho\frac{1-\sum\limits_{m=0}^{\infty}\frac{(-\omega)^{m}}{m!}}{\omega}=\rho\sum\limits_{m=1}^{\infty}\frac{(-\omega)^{m-1}}{m!}.$
The first term in the sum is one, and all other terms converge to zero as $\omega\rightarrow 0$.
Lemma 4.33.
If the Erlang mixture kernel $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is bounded, it is uniformly integrable, i.e., for all $a\in\mathbb{R}_{>0}$ and for every $\epsilon\in\mathbb{R}_{>0}$, there exists a $\Delta t\in\mathbb{R}_{>0}$ such that if $|t_{1}-t_{0}|<\Delta t$, then,
(64) $\int_{t_{0}}^{t_{1}}\hat{\alpha}(t)\,\mathrm{d}t$ $<\epsilon.$
Proof 4.34.
The uniform integrability follows directly from the boundedness:
(65) $\int_{t_{0}}^{t_{1}}\hat{\alpha}(t)\,\mathrm{d}t$ $\leq\int_{t_{0}}^{t_{1}}K\,\mathrm{d}t=K|t_{1}-t_{0}|<K\Delta t.$
Here, $K$ is the upper bound on $\hat{\alpha}$, and the bound is satisfied for $\Delta t=\epsilon/K$.
Lemma 4.35.
Let $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be an infinite-order Erlang mixture approximation with coefficients defined by (17) and rate parameter $a\in\mathbb{R}_{>0}$. Then, $\hat{\alpha}$ is tight, i.e., for every $\epsilon\in\mathbb{R}_{>0}$, there exists a $t_{0}\in\mathbb{R}_{\geq 0}$ and an $\bar{a}\in\mathbb{R}_{>0}$ such that for all $a>\bar{a}$,
(66) $\int_{t_{0}}^{\infty}\hat{\alpha}(t)<\epsilon.$
Proof 4.36.
First, we use a triangle inequality:
(67) $\int_{t_{0}}^{\infty}\hat{\alpha}(t)\,\mathrm{d}t$ $=\int_{t_{0}}^{\infty}\hat{\alpha}(t)\,\mathrm{d}t-\int_{t_{0}}^{\infty}\alpha(t)\,\mathrm{d}t+\int_{t_{0}}^{\infty}\alpha(t)\,\mathrm{d}t$
$\leq\left|\int_{t_{0}}^{\infty}\hat{\alpha}(t)\,\mathrm{d}t-\int_{t_{0}}^{\infty}\alpha(t)\,\mathrm{d}t\right|+\int_{t_{0}}^{\infty}\alpha(t)\,\mathrm{d}t.$
Next, we choose $t_{0}$ large enough that
(68) $\int_{t_{0}}^{\infty}\alpha(t)\,\mathrm{d}t<\epsilon/2.$
Furthermore, we know from Theorem 4.28 that we can choose $\bar{a}$ large enough that
(69) $\left|\int_{t_{0}}^{\infty}\hat{\alpha}(t)\,\mathrm{d}t-\int_{t_{0}}^{\infty}\alpha(t)\,\mathrm{d}t\right|$ $=\left|1-\int_{0}^{t_{0}}\hat{\alpha}(t)\,\mathrm{d}t-\left(1-\int_{0}^{t_{0}}\alpha(t)\,\mathrm{d}t\right)\right|$
$=\left|\int_{0}^{t_{0}}\hat{\alpha}(t)\,\mathrm{d}t-\int_{0}^{t_{0}}\alpha(t)\,\mathrm{d}t\right|<\epsilon/2$
for $a>\bar{a}$. This concludes the proof.
4.5 Convergence of the integral of the absolute error
In this section, we prove another main result, i.e., that the integral of the absolute error of the Erlang mixture approximation converges to zero as the rate parameter goes to infinity. This result is relevant to proving the relation between the stability analyses in Section 5.2 and the pointwise convergence of the state error in Section 5.3.
Theorem 4.37.
Let $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be an Erlang mixture kernel of infinite order with coefficients defined by (17) and rate parameter $a\in\mathbb{R}_{>0}$, and let $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be a regular kernel. Then, for every $\epsilon\in\mathbb{R}_{>0}$, there exists an $\bar{a}\in\mathbb{R}_{>0}$ such that
(70) $\int_{0}^{\infty}|\hat{\alpha}(t)-\alpha(t)|\,\mathrm{d}t$ $<\epsilon$
for all $a>\bar{a}$.
Proof 4.38.
As $\alpha$ is regular, it is bounded. According to Lemma 4.29, this means that $\hat{\alpha}$ is also bounded. Consequently, according to Lemma 4.33, $\hat{\alpha}$ is uniformly integrable. Furthermore, according to Lemma 4.35 and Theorem 4.26, $\hat{\alpha}$ is also tight and converges pointwise to $\alpha$. Finally, according to Vitali’s convergence lemma [48, Sec. 19.1, p. 399], uniform integrability, tightness and pointwise convergence implies that $\hat{\alpha}$ also converges in the $L^{1}$-norm, i.e., the stated result.
@@ 5 Approximate system of ordinary differential equations
In this section, we use the Erlang mixture approximation presented in Section 4 to approximate the multivariate kernel in the DDEs presented in Section 3, and we use the LCT to transform the resulting approximate system to ODEs. Furthermore, we describe the steady state equations for the approximate system of ODEs together with a stability criterion, and we use Hurwitz’ convergence theorem to show that the roots of the characteristic function for the original system of DDEs are accumulation points for the roots of the characteristic function for the approximate system of ODEs. Finally, we analyze the stability of the state error dynamics, and we use the theorem presented by Guglielmi and Hairer [18] to prove that the solution to the approximate system of ODEs converges to that of the original system.
@@
where $\hat{Z}:\mathbb{R}\rightarrow\mathbb{R}^{M+n_{z}}$ contains auxiliary memory states, and $M=\sum_{i=1}^{n_{z}}M_{i}$. Furthermore, the matrices $A\in\mathbb{R}^{M+n_{z}\times M+n_{z}}$, $B\in\mathbb{R}_{\geq 0}^{M+n_{z}\times n_{z}}$, and $C\in[0,1]^{n_{z}\times M+n_{z}}$ are block-diagonal:
@@
5.1 Steady states and stability criteria
In this section, we confirm that the steady state of the approximate system of ODEs satisfies the same equations as that of the original system, and we present the steady state stability criterion. In Section 5.2, we show that the roots involved in the criterion converge to those in the criterion for the original system of DDEs.
@@ 5.1 Steady states and stability criteria
(80) $\left(A^{(i)}\right)^{-1}$ $=\frac{-1}{a_{i}}\begin{bmatrix}1\\ 1&1\\ \vdots&&\ddots\\ 1&1&\cdots&1\end{bmatrix}$, $c^{(i)}\left(A^{(i)}\right)^{-1}b^{(i)}$ $=-\sum_{m=0}^{M_{i}}c_{m,i}=-1.$
Consequently, $CA^{-1}B=I$ is an identity matrix, and
@@
where $F$, $G$, and $H$ are the Jacobian matrices given by (9). Furthermore, the eigenvalues, $\lambda$, that are not equal to $a_{i}$ for $i=1,\ldots,n_{z}$ will also satisfy the reduced characteristic equation
(83) $\hat{P}(\lambda)=\det(F-\lambda I+GQ(\lambda)H)$ $=0,$ $Q_{ij}(\lambda)$ $=\begin{cases}\sum\limits_{m=0}^{M_{i}}c_{m,i}\left(\frac{a_{i}}{a_{i}+\lambda}\right)^{m+1},&i=j,\\ 0,&i\neq j,\end{cases}$
where $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ is the approximate characteristic function and $Q:\mathbb{C}\rightarrow\mathbb{C}^{n_{z}\times n_{z}}$ is an auxiliary matrix.
@@
where we have used Schur’s determinant formula (see, e.g., [57]), which presumes that $\det(A-\lambda I)\neq 0$. Next, we exploit the block-diagonal structure of $A$:
@@
which is nonzero by assumption such that $A-\lambda I$ is invertible. Next, we consider the diagonal matrix
@@
Corollary 5.6.
Let each element of $\hat{\alpha}:\mathbb{R}\rightarrow\mathbb{R}^{n_{z}}$ be an Erlang mixture kernel. Then, the approximate characteristic function $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ and the auxiliary matrix $Q:\mathbb{C}\rightarrow\mathbb{C}^{n_{z}\times n_{z}}$ defined in (83) in Theorem 5.4 are equal to
(92) $\hat{P}(\lambda)$ $=\det(F-\lambda I+G\int_{0}^{\infty}e^{-\lambda s}\hat{\alpha}(s)\,\mathrm{d}sH),$ $Q_{ii}(\lambda)$ $=\int_{0}^{\infty}e^{-\lambda s}\hat{\alpha}_{i}(s)\,\mathrm{d}s,$
for $i=1,\ldots,n_{z}$.
Proof 5.7.
The result follows directly from the equivalence of the approximate systems of DDEs (71) and ODEs (73) and the stability criterion in Lemma 3.5. However, for completeness, we write out the proof in detail. First, we write out the following integral for an Erlang kernel of order $m$ with rate parameter $a_{i}$:
@@
and the characteristic function in (8) becomes
(96) $\hat{P}(\lambda)=\det(F-\lambda I+GQ(\lambda)H).$
@@
Proof 5.9.
As the rate parameters, $a_{i}$, are positive for $i=1,\ldots,n_{z}$, an eigenvalue of zero must be a root of the reduced characteristic function in (83), i.e., it is not an eigenvalue of $A$. Furthermore,